OpenStudy (anonymous):

solve 1/4-5/6=1/b help

OpenStudy (anonymous):

-12/7

OpenStudy (anonymous):

what you have to do is to substitute the operations. get the "b" that is dividing 1 multiplying to the other side, we get (1/4-5/6)b=1

OpenStudy (anonymous):

okay what does that mean @Onlymaths

OpenStudy (anonymous):

The value of b is -12/7.

OpenStudy (skullpatrol):

$$\Huge \frac{1}{4}-\frac{5}{6}=\frac{1}{b}$$

OpenStudy (anonymous):

after you've done that what you have to do is pass the (1/4-5/6) dividing the one; b(1/4-5/6)=1, b= 1/(1/4-5/6) and I believe that you can solve that operation now, cant you?

OpenStudy (skullpatrol):

Do you know how to subtract fractions @gianna?